# Evaluating the Impact of Inter-cluster Communications in Edge Computing

### v2

@@ Abstract
Distributed applications based on micro-services in edge computing are becoming increasingly popular due to the rapid evolution of mobile networks. While Kubernetes is the default framework when it comes to orchestrating and managing micro-service-based applications in mobile networks, the requirement to run applications between multiple sites at cloud and edge poses new challenges. Since Kubernetes does not natively provide tools to abstract inter-cluster communications at the application level, inter-cluster communication in edge computing is becoming increasingly critical to the application performance. In this paper, we evaluate for the first time the impact of inter-cluster communication on edge computing performance by using three prominent, open source inter-cluster communication projects and tools, i.e., Submariner, ClusterLink and Skupper. We develop a fully open-source testbed that integrates these tools in a modular fashion, and experimentally benchmark sample applications, including the ML class of applications, on their performance running in the multi-cluster edge computing system under varying networking conditions. We experimentally analyze two classes of envisioned mobile applications, i.e., a) industrial automation, b) vehicle decision drive assist. Our results show that ClusterLink performs best out of the three tools in scenarios with increased payloads, regardless of the underlying networking conditions or transmission direction between clusters. It is closely followed by Skupper, unless request and reply both transport significant amounts of data. Finally, when requesting smaller amounts of data from a service, Submariner slightly outperforms Skupper and ClusterLink regardless of the inter-node networking conditions.
@@ I Introduction
In this work, we experimentally study and comparatively analyze for the first time three notable open source projects and tools for inter-cluster communication in edge networking and computing: Submariner, ClusterLink and Skupper. While there are several other tools for general inter-cluster communication, this study focuses on open-source solution that do not necessitate sidecar containers or a specific container network interface in the connected clusters, and as such allow for the highest degree of compatibility with existing infrastructure. To this end, we engineer an open-source testbed for inter-cluster communication tools. Its realization uses only production-grade open-source tools, while it provides the ability to vary the underlying network conditions to mimic an edge network infrastructure setting with various link qualities. The inter-cluster communication tools are implemented in a pluggable fashion, thus allowing for a comparative analysis. We also design a set of innovative case studies for novel distributed mobile applications, including use cases based on machine learning (ML), which necessitate inter-cluster communication. Finally, we provide measurements and evaluate the impact on application performance. In the measurements, we attempt to answer the following three research questions, which remain unaddressed by existing work for the scope of inter-cluster communication: 1. How does the distance between clusters impact the service response time of latency sensitive applications? 2. What is the combined effect of cluster distance and payload size on the application’s performance? 3. How do various inter-cluster communication tools comparatively perform for latency sensitive applications?
@@ II-A Related work
Kubernetes is the de facto default tool for managing workloads in the cloud. It organizes computing resources in so-called clusters that can be formed based on different characteristics of nodes, administrative or management goals, or yet a combination of them. Dynamic clusters are formed based on aspects subject to continuous changes, like node proximity in scenarios with moving computing nodes [1]. Static clusters are usually based on hardware properties to form specialized groups, for example, to process tasks that demand hardware accelerators such as GPUs [4]. Such specialized clusters are employed to guarantee the compatibility of applications that demand specific resources since those devices might leverage enhancements to utilize specific resources like GPUs[5]. Additionally, scheduling and workload sharing can differ in those particular installations, which could lead to incompatibility with other Kubernetes clusters [6]. General clustering approaches are analyzed in [3], showing how dependent the selection objectives can be on the specific use case.
@@
Distributed application and service function chains can be formed across multiple clusters. The applications are scheduled and scaled based on node mobility or network properties to minimize service latency[2, 7], increase cost-efficiency, reduce power consumption or boost robustness [3]. Multi-cluster deployments of applications are gaining attention because they address a variety of issues faced in distributed single-cluster environments. On the one hand, there are business reasons to rely on multiple cloud providers (i.e., , multi-cluster environments), such as preventing vendor lock-in and ensuring a global presence. On the other hand, there are technical reasons, such as the need for edge computing to fulfill stringent latency requirements of novel applications (e.g., extended reality) [8], creating scalability issues. In  [9], the authors show that the Kubernetes control plane can become a bottleneck when a certain cluster size is exceeded. This restriction limits the size of clusters since consensus has to be maintained across all member nodes. While the issue can be reduced by changing or improving the used key-value storage system, this does not provide a complete solution, and the problem can be expect to increase relative to the geographical spread of the nodes. Additionally, due to the inherent characteristics of edge nodes (i.e., constrained computing and storage capacity), some applications can only partially be executed in an edge cluster, which in turn requires communication with other clusters.
@@
In general, research focusing on the inter-cluster communication problem to allow multi-cluster collaboration is work in progress. In [10], the authors perform an assessment of Submariner’s performance as well as the distributed application deployed on top of it with a focus on CPU utilization. However, this study lacks consideration of two critical dimensions: varying inter-node/inter-cluster latency and the amount of data being exchanged between the clusters. In [11] the authors explore inter-cluster communication by proposing a framework to deploy a 5G experimental network across distributed infrastructure formed by multiple geographically spread clusters. The authors evaluated the maximum achievable throughput in inter-cluster communications provided by Submariner using three different tunneling options (i.e., Wireguard, Libreswan, and VXLAN) and three different transport protocols (TCP, UDP and SCTP). These works focus on Submariner while our focus is on a comparative analysis in various scenarios with ClusterLink and Skupper.
@@
Focusing on cloud-driven scenarios for mobile networks furthermore, paper [12] proposed a solution for multi-cluster management and connectivity employing two open-source tools: KubeFed and NSM. The Libreswan technology was used to tunnel traffic between clusters. The approach proposed utilized TCP, UDP, and SCTP as transport protocols and benchmarked basic data plane performance, collecting maximum throughput and latency metrics between containers. Since latency between nodes was static, real-world performance in the network edge would diverge significantly from these findings. A valuable extension of paper [12] would in fact be to compare the tool proposed with some of the production-grade tools that we use in our paper.
@@
Related work in benchmarking of intra-cluster networking solutions, such as in [13] focuses on total performance within a single cloud cluster, which varies based on the hardware used. Paper [14] conducted an extensive survey of benchmarking performance at the edge and observed that further research is needed to measure the performance of edge networks efficiently. This survey reveals a gap in explicitly benchmarking orchestrators and other software that connects the infrastructure into an edge continuum. Indeed, many benchmarks are limited by physical aspects of edge devices (e.g., processing power, chipset throughput, storage, and memory). Our evaluation differs from previous works as, instead of absolute metrics like resource utilization, it focuses on the relative response time differences encountered through varying aspects of 1) underlying network, 2) payload size, and 3) direction of payload exchange between clusters.
@@ II-B Inter-cluster Communication Tools
Tab. I lists inter-cluster communication tools that we implement in this paper; Submariner [^1] https://submariner.io/ (CNCF, RedHat) and Skupper [^2] https://skupper.io/, along with the emerging ClusterLink [^3] https://clusterlink.net/ (IBM) [15]. While all three of them tackle the same challenge that is of inter-cluster service communication, each tool has unique features that set it apart. Skupper, being a layer 7 interconnect, does not need any further configuration of the CIDR ranges that are used to assign IP addresses to pods or services, making it scale very well with comparably low configuration effort. Submariner on the other hand works on layer 3 and employs a broker for connection establishment, leading to only this one component having to be publicly accessible; it can potentially even be hosted on hardware that is not part of the communicating clusters. This allows for cluster-to-cluster communication between dynamic endpoints through NAT as well as firewalls. Using HTTP, ClusterLink shares the benefits of layer 7 connections but also provides authentication as well as granular permission and access management on workload, service and cluster level, allowing for a high degree of security and low levels of implicit trust between endpoints. All three solutions are written in Go, provide state-of-the-art encryption out of the box and are deployed on top of the respective Kubernetes clusters without the need for modifications of the individual node’s operating systems.
@@ II-C Applications
The performance evaluation of the inter-cluster communication tools implemented uses key performance indicators (KPIs) and characteristics of the emerging mobile computing applications. We choose case studies of two classes of applications: industrial automation and cellular Vehicle-to-Everything (C-V2X). They are illustrative of the real-world verticals that require robust inter-cluster communication, i.e.,
@@ III System Architecture and Testbed Design
The software stack used consists of hierarchical layers, as seen in Fig. 1. Since we publish our code base on Github [^4] https://github.com/michalke-it/inter-cluster-benchmarking, we design the testbed exclusively around open-source tools, akin to [22]. The functionalities provided by each layer are described below. This environment guarantees a common baseline for all three tools. Most notably, our testbed allows for the manipulation of network properties like delay, jitter, and packet loss through software, while keeping all other factors static, thus enabling comparative performance measurements.
@@ III-C Applications
The three example applications are functions served on top of the Knative [^5] https://knative.dev framework which handles requests and provides HTTP endpoints. Their source code is published on Github [^6] https://github.com/fcarp10/serverless-functions. While all of these applications are deployed at the same time, they are interacted with sequentially. Automatic scaling in Kubernetes is deactivated since we do not aim to measure total achievable performance on the given hardware but instead relative response time differences. The functionality of the respective functions can be summarized as follows:
@@
This function receives any payload sent as the HTTP body and acknowledges it with an HTTP 200 response without body. It is used for the Remote Maintenance (ML Training) application, since we can see the same data transfer here as in the previous function with the difference that the data is not transferred back. This behavior would be expected whenever an ML model is trained, since the sending cluster would not have to wait for the training to complete.
@@ IV Evaluation Setup
The main challenge in the evaluation setup is to create the testing environment and applications as close as possible to the real world use case so that meaningful results for the application response times can be obtained. To this end, and based on the description provided in Section II, we map the analyzed communication patterns to our example applications, and design the matching Jmeter tests to evaluate them. We then gather the response times achieved through the different tools across a set of six different infrastructure scenarios that mimic edge environments with different characteristics. To create a baseline, we start with a test without any interface manipulation, leaving only the delay, jitter and packet loss experienced physically by the transmission. This establishes a reference value for each tool when running on our provided hardware and software stack.
We then start introducing delay, jitter and packet loss to the frames transmitted through the Ethernet interface, emulating a transmission over higher distances or lower quality connection media between all nodes as listed in Tab. IV. The naming of each parameter set is based on the additional delay introduced to an average round trip time between any two nodes; e.g. ,10ms in edg10. We base our values on the measurements described by the authors in [8] and extrapolate them to gain a set of test values.
@@ IV-A Applications
All tests are conducted based on Fig. 2, including the request and response direction. Tab. V details the payload size parameters of the files used to generate the requests/responses.
@@ IV-A1 Remote maintenance inference
Use case Payload size
@@ IV-A2 Remote Maintenance (ML Training)
In this test, using the function payload-recv, we emulate the behavior of the system sending data for training an ML model. The data is assumed to have been aggregated, resulting in variable-sized data in a cluster that is sent to a second cluster. Since the data is used for ML training, the destination cluster, cluster 1 in Fig. 2, only acknowledges receiving the data without any payload. The payload of the request ranges from 60B to 1MB, while the payload of the response is zero bytes (0B) as detailed in Table V. Since machine learning training has high tolerance regarding the response time, depending on the application, we choose an arbitrary maximum response time of 1000 ms. This allows us to compare the effect of a simple acknowledgement to the extensive application response of the previous test.
@@ IV-A3 Vehicle decision assist
Here we use the static-reply function. A car in a platoon is overtaking and coordinating with either the roadside units of the section, or alternatively/complementary an oncoming platoon of vehicles. It has to receive data from that other cluster, e.g., platoon length, speed, and position, with low tolerance for delay. Therefore, we assume an empty payload for the request, which aims to trigger a response with a fixed payload size of 1000B, as shown in Table V.
@@ V-A Remote Maintenance (ML Inference)
For smaller payloads, ClusterLink and Submariner performed similarly until the eg10 value set, with Submariner being slightly outperformed by ClusterLink in most cases as depicted in Fig. 4. However, the difference is subtle and in consideration of run to run variance, their performance can be considered equal here. Skupper, while initially delivering comparable performance to the other tools breaks away in the edg15 scenario by delivering ever higher response times, by almost 20% in the edg20 scenario, indicating a downside of using AMQP in this environment. From a response time perspective, this makes Skupper the last choice out of the three tools for inter-cluster communication when dealing with applications that need to send medium-sized payloads to a service that in turn replies with small payloads.
@@
We can therefore conclude that for workloads where small amounts of data are sent to remote services that reply with a similar amounts of data, the three tools perform similarly to each other. For medium and larger data, i.e. 100 KB and more, the tool selection becomes a decision of significant impact.
@@ V-B Remote Maintenance (ML Training)
Fig. 6 shows the results of the ML training case studies for 1MB payload size, which, differently from the inference use case, does not consider communication back to the request source besides acknowledgments. Again, we can observe ClusterLink and Skupper behaving similarly to each other with slightly lower medians for ClusterLink, while it now also exhibits a slightly higher spread. Similarly to the ML inference for the remote maintenance use case with payloads of 1 MB, Submariner performs worst among the tested tools. Here, it presents about four times the median values of ClusterLink and Skupper in the edg20 scenario along with a high standard deviation in the observed data. In contrast to the previous use case, Submariner exceeds the boundary of 1000 ms already in the edg10 scenario with some measurements and in the edg15 scenario with its median value. ClusterLink and Skupper however remain below the threshold with their median values but slightly exceed the boundary in the edg20 scenario with a small portion of the observed samples.
@@
Curiously, the worst performance of Submariner that we observed was when data was not sent back to the tester container (or client), apart from acknowledgments. We can therefore conclude, that Submariner’s implementation seems to have suboptimal handling of HTTP POST responses when it comes to larger payloads, which makes it unsuitable for this application when network latency is above 7.5 ms in each direction (assuming a symmetric link, therefore 15 ms for round trip).

### v3

@@ Abstract
Distributed applications based on micro-services in edge computing are becoming increasingly popular due to the rapid evolution of mobile networks. While Kubernetes is the default framework when it comes to orchestrating and managing micro-service-based applications in mobile networks, the requirement to run applications between multiple sites at cloud and edge poses new challenges. Since Kubernetes does not natively provide tools to abstract inter-cluster communications at the application level, inter-cluster communication in edge computing is becoming increasingly critical to the application performance. In this paper, we evaluate for the first time the impact of inter-cluster communication on edge computing performance by using three prominent, open source inter-cluster communication projects and tools, i.e., Submariner, ClusterLink and Skupper. We develop a fully open-source testbed that integrates these tools in a modular fashion, and experimentally benchmark sample applications, two of them centered around ML, on their performance running in a multi-cluster edge computing system under varying networking conditions. We experimentally analyze two classes of envisioned mobile applications, i.e., a) industrial automation, b) vehicle decision drive assist. Our results show that ClusterLink performs best out of the three tools in scenarios with increased payloads, regardless of the underlying networking conditions or transmission direction between clusters. It is closely followed by Skupper, unless request and reply both transport significant amounts of data. Finally, when requesting smaller amounts of data from a service, Submariner slightly outperforms Skupper and ClusterLink regardless of the inter-node networking conditions.
@@ I Introduction
In this work, we experimentally study and comparatively analyze for the first time three notable open source projects and tools for inter-cluster communication in edge networking and computing: Submariner, ClusterLink and Skupper. While there are several other tools for general inter-cluster communication, this study focuses on open-source solutions that do not necessitate sidecar containers or a specific container network interface in the connected clusters, and as such allow for the highest degree of compatibility with existing infrastructure. To this end, we engineer an open-source testbed for inter-cluster communication tools. Its realization uses only production-grade open-source tools, while it provides the ability to vary the underlying network conditions to mimic an edge network infrastructure setting with various link qualities. The inter-cluster communication tools are implemented in a pluggable fashion, thus allowing for a comparative analysis. We also design a set of innovative case studies for novel distributed mobile applications, including use cases based on machine learning (ML), which necessitate inter-cluster communication. Finally, we provide measurements and evaluate the impact on application performance. In the measurements, we attempt to answer the following three research questions, which remain unaddressed by existing work for the scope of inter-cluster communication: 1. How does the distance between clusters impact the service response time of latency sensitive applications? 2. What is the combined effect of cluster distance and payload size on the application’s performance? 3. How do various inter-cluster communication tools comparatively perform for latency sensitive applications?
@@ II-A Related work
Kubernetes is the de facto default tool for managing workloads in the cloud. It organizes computing resources in so-called clusters that can be formed based on different characteristics of nodes, administrative or management goals, or yet a combination of them. Dynamic clusters are formed based on aspects subject to continuous changes, like node proximity in scenarios with moving compute nodes [1]. Static clusters are usually based on hardware properties to form specialized groups, for example, to process tasks that demand hardware accelerators such as GPUs [4]. Such specialized clusters are employed to guarantee the compatibility of applications that demand specific resources since those devices might leverage enhancements to utilize specific resources like GPUs[5]. Additionally, scheduling and workload sharing can differ in those particular installations, which could lead to incompatibility with other Kubernetes clusters [6]. General clustering approaches are analyzed in [3], showing how dependent the selection objectives can be on the specific use case.
@@
Distributed application and service function chains can be formed across multiple clusters. The applications are scheduled and scaled based on node mobility or network properties to minimize service latency[2, 7], increase cost-efficiency, reduce power consumption or boost robustness [3]. Multi-cluster deployments of applications are gaining attention because they address a variety of issues faced in distributed single-cluster environments. On the one hand, there are business reasons to rely on multiple cloud providers (i.e., , multi-cluster environments), such as preventing vendor lock-in and ensuring a global presence. On the other hand, there are technical reasons, such as the need for edge computing to fulfill stringent latency requirements of novel applications (e.g., extended reality) [8], creating scalability issues. In  [9], the authors show that the Kubernetes control plane can become a bottleneck when a certain cluster size is exceeded. This restriction limits the size of clusters since consensus has to be maintained across all member nodes. While the issue can be reduced by changing or improving the used key-value storage system, this does not provide a complete solution, and the problem can be expected to increase relative to the geographical spread of the nodes. Additionally, due to the inherent characteristics of edge nodes (i.e., constrained computing and storage capacity), some applications can only partially be executed in an edge cluster, which in turn requires communication with other clusters.
@@
In general, research focusing on the inter-cluster communication problem to allow multi-cluster collaboration is work in progress. In [10], the authors perform an assessment of Submariner’s performance as well as the distributed application deployed on top of it with a focus on CPU utilization. However, this study lacks consideration of two critical dimensions: varying inter-node/inter-cluster latency and the amount of data being exchanged between the clusters. In [11] the authors explore inter-cluster communication by proposing a framework to deploy a 5G experimental network across distributed infrastructure formed by multiple geographically spread clusters. The authors evaluated the maximum achievable throughput in inter-cluster communications provided by Submariner using three different tunneling options (i.e., Wireguard, Libreswan, and VXLAN) and three different transport protocols (TCP, UDP and SCTP). While the cited works focus solely on Submariner, we compare the three main tools used for inter-cluster communication in multiple scenarios.
@@
Focusing on cloud-driven scenarios for mobile networks furthermore, paper [12] proposed a solution for multi-cluster management and connectivity employing two open-source tools: KubeFed and NSM. The Libreswan technology was used to tunnel traffic between clusters. The proposed approach employs TCP, UDP, and SCTP as transport protocols, and the authors benchmark basic data plane performance, collecting maximum throughput and latency metrics between containers. Since latency between nodes was static, real-world performance in the network edge would diverge significantly from these findings. A valuable extension of paper [12] would in fact be to compare the tool proposed with some of the production-grade tools that we use in our paper.
@@
Related work in benchmarking of intra-cluster networking solutions, such as in [13], focuses on total performance within a single cloud cluster, which is hardware-dependent. Paper [14] conducted an extensive survey of benchmarking performance at the edge and observed that further research is needed to measure the performance of edge networks efficiently. This survey reveals a gap in explicitly benchmarking orchestrators and other software that connects the infrastructure into an edge continuum. Indeed, many benchmarks are limited by physical aspects of edge devices (e.g., processing power, chipset throughput, storage, and memory). Our evaluation differs from previous works as it focuses on the relative response time differences encountered through varying aspects of 1) underlying network, 2) payload size, and 3) direction of payload exchange between clusters, instead of absolute metrics like resource utilization.
@@ II-B Inter-cluster Communication Tools
Table I lists the inter-cluster communication tools that we implement in this paper: Submariner [^1] https://submariner.io/ (CNCF, RedHat) and Skupper [^2] https://skupper.io/, along with the emerging ClusterLink [^3] https://clusterlink.net/ (IBM) [15]. While all three tools tackle the same challenge of inter-cluster service communication, each has unique features that set them apart. Skupper, being a layer 7 interconnect, does not need any further configuration of the CIDR ranges that are used to assign IP addresses to pods or services, making it scale very well with comparably low configuration effort. Submariner on the other hand works on layer 3 and employs a broker for connection establishment, leading to only this one component having to be publicly accessible; it can potentially even be hosted on hardware that is not part of the communicating clusters. This allows for cluster-to-cluster communication between dynamic endpoints through NAT as well as firewalls. Using HTTP, ClusterLink shares the benefits of layer 7 connections but also provides authentication as well as granular permission and access management on workload, service and cluster level, allowing for a high degree of security and low levels of implicit trust between endpoints. All three solutions are written in Go, provide state-of-the-art encryption out of the box and are deployed on top of the respective Kubernetes clusters without the need for modifications of the individual node’s operating systems.
@@ II-C Applications
The performance evaluation of the inter-cluster communication tools implemented uses key performance indicators (KPIs) and characteristics of emerging mobile computing applications. We choose case studies of two classes of applications: industrial automation and cellular Vehicle-to-Everything (C-V2X). These are illustrative of the real-world verticals that require robust inter-cluster communication and are described as follows.
@@ III System Architecture and Testbed Design
The software stack used consists of hierarchical layers, as seen in Fig. 1. Since we publish our code base on Github[22], we design the testbed exclusively around open-source tools, akin to [23]. The functionalities provided by each layer are described below. This environment guarantees a common baseline for all three tools. Most notably, our testbed allows for the manipulation of network properties like delay, jitter, and packet loss through software, while keeping all other factors static, thus enabling comparative performance measurements.
@@ III-C Applications
The three example applications are functions served on top of the Knative[24] framework, which handles requests and provides HTTP endpoints. Their source code is published on GitHub[25]. While all of those applications are deployed at the same time, we interact with them sequentially. Automatic scaling in Kubernetes is deactivated since we do not aim to measure total achievable performance on the given hardware but instead relative response time differences. The functionality of each deployed function can be described as follows:
@@
This function receives any payload sent as HTTP body and acknowledges it with an HTTP 200 response without body. It is used for the Remote Maintenance (ML Training) application, since we can see the same data transfer here as in the previous function with the difference that the data is not transferred back. This behavior would be expected whenever an ML model is trained, since the sending cluster would not have to wait for the training to complete.
@@ IV Evaluation Setup
The main challenge in the evaluation setup is to create the testing environment and applications as close as possible to the real world use case such that meaningful results for the application response times can be obtained. To this end, and based on the description provided in Section II, we map the analyzed communication patterns to our example applications, and design the matching Jmeter tests to evaluate them. We then gather the response times achieved through the different tools across a set of six different infrastructure scenarios that mimic edge environments with different characteristics. To create a baseline, we start with a test without any interface manipulation, leaving only the delay, jitter and packet loss experienced physically by the transmission. This establishes a reference value for each tool when running on our provided hardware and software stack.
We then start introducing delay, jitter and packet loss to the frames transmitted through the Ethernet interface, emulating a transmission over higher distances or lower quality connection media between all nodes as listed in Table IV. The naming of each parameter set is based on the additional delay introduced to an average round trip time between any two nodes; e.g., 10ms in edg10. We base our values on the measurements described by the authors in [8] and extrapolate them to gain a set of test values.
@@ IV-A Applications
All tests are conducted based on Fig. 2, including the request and response direction. Table V details the payload sizes of the files used to generate the requests/responses.
@@ IV-A1 Remote maintenance inference
Use case HTTP Payload size
@@ IV-A2 Remote maintenance (ML training)
In this test, using the function payload-recv, we emulate the behavior of the system sending data for training an ML model. The data is assumed to have been aggregated, resulting in variable-sized data in a cluster that is sent to a second cluster. Since the data is used for ML training, the destination cluster, cluster 1 in Fig. 2, only acknowledges receiving the data without sending any payload. The payload of the request ranges from 60B to 1MB, while the payload of the response is zero bytes (0B) as detailed in Table V. Since, depending on the application, machine learning training can have high tolerance regarding the response time we choose an arbitrary maximum response time of 1000 ms. This allows us to compare the effect of a simple acknowledgement to the extensive application response of the previous test.
@@ IV-A3 Vehicle decision assist
Here, we use the static-reply function. A car in a platoon is overtaking and coordinating with either the roadside units of the section, or alternatively/complementary an oncoming platoon of vehicles. It has to receive data from the other cluster, e.g., platoon length, speed, and position, with low tolerance for delay. Therefore, we assume an empty payload for the request, which aims to trigger a response with a fixed payload size of 1000B, as shown in Table V.
@@ V-A Remote Maintenance (ML Inference)
For smaller payloads, ClusterLink and Submariner perform similarly until the eg10 value set, with Submariner being slightly outperformed by ClusterLink in most cases as depicted in Fig. 4. However, the difference is subtle and in consideration of run to run variance, their performance can be considered equal here. Skupper, while initially delivering comparable performance to the other tools breaks away in the edg15 scenario by delivering even higher response times, by almost 20% in the edg20 scenario, indicating a downside of using AMQP in this environment. From a response time perspective, this makes Skupper the last choice out of the three tools for inter-cluster communication when dealing with applications that need to send medium-sized payloads to a service that in turn replies with small payloads.
@@
We can therefore conclude that for workloads where small amounts of data are sent to remote services that reply with similar amounts of data, the three tools perform similarly to each other. For medium and larger data, i.e. 100 KB and more, the tool selection becomes a decision of significant impact.
@@ V-B Remote Maintenance (ML Training)
Fig. 6 shows the results of the ML training case studies for 1MB payload size, which, differently from the inference use case, does not consider communication back to the request source besides acknowledgments. We can observe ClusterLink and Skupper behaving similarly to each other with slightly lower medians for ClusterLink, while it now also exhibits a slightly higher spread. Similarly to the ML inference for the remote maintenance use case with payloads of 1 MB, Submariner performs worst among the tested tools. Here, it presents about four times the median values of ClusterLink and Skupper in the edg20 scenario along with a high standard deviation in the observed data. In contrast to the previous use case, Submariner exceeds the boundary of 1000 ms already in the edg10 scenario with some measurements and in the edg15 scenario with its median value. ClusterLink and Skupper however remain below the threshold with their median values but slightly exceed the boundary in the edg20 scenario with a small portion of the observed samples.
@@
Curiously, we observe the worst performance of Submariner when data is not sent back to the tester container (or client), apart from acknowledgments. We can therefore conclude, that Submariner’s implementation seems to have suboptimal handling of HTTP POST responses when it comes to larger payloads, which makes it unsuitable for this application when network latency is above 7.5 ms in each direction (assuming a symmetric link, therefore 15 ms for round trip).